OpenStudy (anonymous):

Find −3a2 − b3 + 3c2 − 2b3 if a = 2, b = −1, and c = 3.

OpenStudy (anonymous):

-3(2)^2 +1^3 + 3(3)^2 – 2(-1)^3

Vocaloid (vocaloid):

^ ok, you've plugged in the numbers correctly, all you have to do is calculate what that expression is equal to...

OpenStudy (anonymous):

I got 14 but I dont know if its correct

Vocaloid (vocaloid):

well, let's take it step by step: -3(2)^2 = ?

OpenStudy (anonymous):

-6^2 or -3(4)?

Vocaloid (vocaloid):

it's -3(4) remember, the exponent only applies to the number next to it

OpenStudy (anonymous):

Okay -3(4)

Vocaloid (vocaloid):

1^3 = ?

OpenStudy (anonymous):

1

Vocaloid (vocaloid):

good. 3(3)^2 = ? (again, remember that the exponent only applies to the number next to it)

OpenStudy (anonymous):

3(9)

Vocaloid (vocaloid):

good, – 2(-1)^3 = ?

OpenStudy (anonymous):

Above, why would the 1 be positive if it was negative in the equation?

OpenStudy (anonymous):

-2(-1)

Vocaloid (vocaloid):

if b = -1 then -b^3 = -(-1)(-1)(-1) = 1

OpenStudy (anonymous):

Ohhhhhh. Okay I get it. Gotcha. Okay continue (:

Vocaloid (vocaloid):

and yes, – 2(-1)^3 = -1 so we put everything together: -3(2)^2 +1^3 + 3(3)^2 – 2(-1)^3 = -3(4) + 1 + 3(9) -2(-1) can you take over from here?

Vocaloid (vocaloid):

g2g but the final answer should be 18

OpenStudy (anonymous):

Got it. Thank you so much